# In-Context Translation: Towards Unifying Image Recognition, Processing, and Generation

### v1

@@ Abstract
We propose In-Context Translation (ICT), a general learning framework to unify visual recognition (e.g., semantic segmentation), low-level image processing (e.g., denoising), and conditional image generation (e.g., edge-to-image synthesis). Thanks to unification, ICT significantly reduces the inherent inductive bias that comes with designing models for specific tasks, and it maximizes mutual enhancement across similar tasks. However, the unification across a large number of tasks is non-trivial due to various data formats and training pipelines. To this end, ICT introduces two designs. Firstly, it standardizes input-output data of different tasks into RGB image pairs, e.g., semantic segmentation data pairs an RGB image with its segmentation mask in the same RGB format. This turns different tasks into a general translation task between two RGB images. Secondly, it standardizes the training of different tasks into a general in-context learning, where “in-context” means the input comprises an example input-output pair of the target task and a query image. The learning objective is to generate the “missing” data paired with the query. The implicit translation process is thus between the query and the generated image. In experiments, ICT unifies ten vision tasks and showcases impressive performance on their respective benchmarks. Notably, compared to its competitors, e.g., Painter and PromptDiffusion, ICT trained on only 4$\times$×RTX 3090 GPUs is shown to be more efficient and less costly in training.
Keywords:
In-Context learning Image translation Unified models
1 Introduction
@@
In the computer vision (CV) community, a unified framework for different tasks is also a long-standing aspiration. The unification is appealing because it side-steps task-specific designs, minimizes the inductive bias inherent in crafting models for different tasks, and maximizes the potential of mutual enhancement among related tasks. However, the progress of such unification in CV lags much behind NLP, due to two main Challenges. C1: NLP has the language token as an intuitive and unified data format for all tasks. However, CV encompasses highly heterogeneous data formats, e.g., images, semantic segmentation masks, keypoints, depth maps, etc. Such variation impedes the unification of expert models of different CV tasks. C2: Most NLP tasks can be easily converted into a sequence completion task [7] as language itself is sequential. In CV, the popular solution for image completion is Masked Autoencoders (MAE) [24]. The objective of MAE is to acquire an image representation by reconstructing the original image from a version where a random subset of image patches is masked out. This representation can be tailored to various specific discrimination tasks through fine-tuning. However, it’s not intuitive how it can be applied to unify other categories of vision tasks, such as conditional image generation.
@@
Thanks to ICT’s unifying designs, in practice, it can be deployed to train various models according to the degree of unification. 1) When trained on joint data of different categories of tasks (e.g., semantic segmentation, image denoising, and depth-to-image generation), it yields a compact model that can generate outputs for every trained task. 2) When trained on joint data of a specific category (e.g., mask-to-image and depth-to-image data for the category of conditional image generation), it derives a single-category model. 3) When trained on data of an individual task, it learns a dedicated single-task model. This is a special case where the model itself is not unified. We highlight that the above three models are trained in different regimes and datasets, yielding distinct model parameters, but the training framework (including input-output format, backbone, training method, etc.) always remains the same, underscoring the “unifying” essence.
@@
Our experiments showcase extensive results of ICT in three types consisting of ten tasks in total. Our key observation is that ICT exhibits promising performance on all categories of tasks, while its competitors Painter [66] and PromptDiffusion [69] collapse on at least one of them (Fig. 1). Our contribution is thus three-fold: 1) a general learning framework that can cope with visual recognition, low-level image processing, and conditional image generation tasks; 2) a new in-context translation method that can be built on an SD model, adapting SD’s pre-trained knowledge for downstream tasks in a unified fashion; and 3) extensive experiments on ten vision tasks and for three types of model training, by which we hope to spur more interesting research on how to induce a profound understanding of vision tasks through a unified learning framework.
2 Related Works
Unified Vision Models. Encouraged by the success of language generalist models [7, 12, 64], training a generalist model to unify different computer vision tasks has attracted significant interest in recent years. We identify three key components of designing a unified vision model: 1) data format, 2) backbone, and 3) training pipeline. Regarding these, we discuss related works and their primary differences from ours.
Some attempts [65, 9, 29, 10, 39] map the input image into discrete representations and implement prompt learning in the discrete space to handle different tasks. A representative work, Unified IO [39], homogenizes various vision data modalities into a sequence of discrete vocabulary tokens and utilizes VQGAN [18] to support dense prediction tasks. This unification of data format via the discretization process leads to lossy data compression, making the solution suboptimal. In contrast, our unified RGB image format better suits vision tasks.
@@ 2 Related Works
Similar to our ICT, a bunch of works [4, 66, 69, 23, 21] utilize the RGB image as a general data format to unify vision tasks. MAE-VQGAN [4] and Painter [66] propose a masked image generation framework, where the input image and an example pair are stitched together and the model needs to predict the masked region (similar to MAE [24]). They adopt a pre-trained ViT-based MAE as the backbone, limiting their performance in image generation. ICT does not suffer from this issue due to the generative backbone we use. In addition, they apply random masking of image patches during training while ICT is trained to predict the whole desired output image that is more in line with the inference scenario. ICT supports the usage of textual prompt to enable more controlled image generation, which is beyond the reach of MAE-VQGAN and Painter.
@@
InstructDiffusion [23] and InstructCV [21] learn a unified model based on a generative backbone and cast different vision tasks as text-guided image editing. Nevertheless, they heavily rely on delicate prompt design, limiting their applications to vision tasks that are difficult to instruct by using human language (e.g., InstructDiffusion cannot handle the task of depth estimation). ICT refrains from this issue by applying a simple image-based context to inform the desired task.
Diffusion Models. Diffusion models [58, 26, 59, 62, 61] have recently become the primary choices for generative modeling of data. Following the definition in Denoising Diffusion Probabilistic Models (DDPM) [26], a diffusion model consists of a forward diffusion process that gradually adds noise to data and a reverse denoising process that reconstructs desired data from the noise. Recent methods based on diffusion models achieve state-of-the-art results in many vision tasks, including image/video generation [16, 41], image editing [42, 6], low-level image processing [54, 67], etc. Notably, large-scale text-to-image diffusion models show compelling ability [45, 53, 50, 3] to generate high-fidelity visual content. These pre-trained models are broadly utilized in applications such as concept customization [20, 52] and image composition [40]. Some recent works [30, 14, 83, 75, 63] further reveal the potential of applying pre-trained diffusion models to discriminative tasks, which encourages us to build a general framework that unifies generative and discriminative model training.
In-Context Learning. GPT-3 [7] has demonstrated the ability of pre-trained LLMs to adapt to various downstream NLP tasks given example pairs. After that, there have been efforts in exploring the ability of in-context learning [38, 43, 51, 44, 71, 55]. A prominent track of works [68, 46, 13, 73] fine-tunes language models on a collection of datasets described via instructions (e.g., task prompt, demonstration examples, and constraints), and attains significant improvement of the model’s generalization on unseen tasks. This in-context learning pipeline has recently been introduced to vision-language tasks as well [2, 37, 22, 31]. A representative work, Flamingo [2], bridges pre-trained vision and language models by fine-tuning them on text-image instruction-following data and showcases impressive few-shot results in a variety of tasks such as image captioning and visual question-answering. By comparison, ICT exploits a new approach of in-context translation, which is based on the use of an image-label pair as well as an optional prompt for both image- and text-level context.
3 In-Context Translation (ICT)
@@
To learn function $f$, we first construct context-based training data by converting the input-output data into RGB images consistently across different vision tasks (Sec. 3.1). Then, we build the ICT framework on top of SD [50] and offer different training strategies according to the unification degrees (Sec. 3.2). For example, for the most unification, each model training iteration uses a batch of context-based data sampled from ten vision tasks (across three types).
3.1 Context-based Data Construction
@@
Visual Recognition. We consider three representative recognition tasks, including semantic segmentation, depth estimation, and keypoint detection. 1) Semantic segmentation is a dense classification task wherein the input is an image and the output is per-pixel semantic labels. We transfer the output to an RGB image by assigning a color to each pixel according to a predefined semantic-color codebook, ensuring that each semantic class has its unique RGB value [66]. During inference, we obtain the predicted semantic class of each pixel by finding its “nearest” color in the codebook. 2) Depth estimation is a dense regression task that needs to predict the depth value of each pixel of the input image. To convert the output, i.e., one-channel depth map, into an RGB image, we linearly scale the range of pixel values to $\left[0,255\right]$ and replicate it three times to yield a three-channel “image”. 3) Keypoint detection aims to locate key object components on the input image. We formulate the output as a RGB image by first drawing squares whose center is the location of keypoints, and then rendering each square with a unique color associated with the key component. ***Due to space limits in the main paper, the hyperparameters for drawing squares, rendering, etc., are given in the supplementary document.
@@
3.2 Training Framework
We implement ICT to run an in-context image translation process on top of SD. SD is a text-to-image model that incorporates a diffusion process in the latent space of a pre-trained autoencoder. Specifically, a denoising U-Net is trained to fit the distribution of latent codes, which models the reverse diffusion process. Taking the noisy latent and the time step as input, this U-Net is further conditioned on the textual embeddings extracted by the text encoder of CLIP [47]. Visual and textual embeddings interact by cross-attention inside the U-Net to produce the output of the current time step. During inference, SD performs iterative reverse diffusion on a randomly sampled noise to generate an image that faithfully adheres to the input text. In our work, to fulfill Eq. 1 for various visual tasks, we fine-tune the U-Net on our context-based datasets.
@@ 3.2 Training Framework
$\mathcal{L}(\theta)=\mathbb{E}_{z,m,b,y,\epsilon\sim\mathcal{N}(0,1),t}\Bigl{[}\lVert\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}(m),b,c_{\phi}(y))\rVert_{2}^{2}\Bigr{]},$ (2)
where $z$ is the latent code extracted from the ground truth, $y$ is the input text, $\epsilon$ is a noise term, $t$ is the time step, $\epsilon_{\theta}$ is the denoising U-Net, $z_{t}$ is the noisy version of $z$ at time $t$, $\mathcal{E}$ is the VAE encoder, and $c_{\phi}$ is the text encoder. We fine-tune the U-Net and keep the text encoder and the autoencoder frozen.
@@
4 Experiments
4.1 Experimental Settings
Table 1: Task and dataset configuration.
Task Dataset Training images Testing images
Segmentation ADE20K 20,210 2,000
Depth estimation NYUv2 24,231 654
Keypoint detection COCO 149,781 6,352
Deraining merged deraining datasets 13,712 4,300
Denoising SIDD 96,000 1,280
Enhancement LoL 485 15
Mask-to-image ADE20K 20,210 2,000
Depth-to-image NYUv2 24,231 654
Pose-to-image COCO 149,781 6,352
Edge-to-image COCO 118,287 5,000
Table 2: Comparison results on visual recognition. : Specialized methods for each task. : Officially trained Painter model using 32$\times$ the computing power of ICT. : Retrained Painter using its official code under the same computing resources as ICT. §: A concurrent work (results are obtained using the official weights). Bold: Best. Underline: Second best. We ignore specialized models when ranking best and second best and this applies to all tables. The results of ICT are reported as the average scores and standard deviations across three trials.
@@ 4.1 Experimental Settings
Table 3: Comparison results on low-level image processing. : Specialized methods for each task. : Officially trained Painter model using 32$\times$ the computing power of ICT. : Retrained Painter using its official code under the same computing resources as ICT. : Following InstructDiffusion [23], it directly reconstructs the ground truth via the autoencoder of pre-trained SD, and the corresponding results indicate the upper bound of ICT. Bold: Best. Underline: Second best.
@@
Table 4: Comparison results on conditional image generation. : Specialized methods for each task. : Trained using official code under the same computing resources as ICT. Note that there is no officially trained Painter model for conditional image generation. Bold: Best. Underline: Second best.
@@
ControlNet [81] 35.4 43.9 43.0 12.9
Painter [66] 75.7 89.3 200.1 233.1
PromptDiffusion [69] 31.0 52.5 40.6 13.8
ICT-st 29.9 $\pm$ 0.3 44.0 $\pm$ 0.7 34.7 $\pm$ 0.3 13.6 $\pm$ 0.2
ICT-sc 27.8 $\pm$ 0.6 44.2 $\pm$ 0.8 34.3 $\pm$ 0.5 13.5 $\pm$ 0.4
Figure 3: Visual comparison results. “GT” represents ground truth. Here we showcase the results on two representative tasks for each category due to space limits. Please see more visualizations in the supplementary.
@@
Datasets. We conduct experiments involving ten vision tasks on six datasets, including ADE20K [84], NYUv2 [57], COCO [36], merged deraining datasets [19, 76, 80, 79, 34], SIDD [1], and LoL [70]. We adopt the same training/testing split as [66]. Please refer to Tab. 1 for a detailed task and dataset configuration.
@@
Methods. We evaluate ICT with its two direct competitors, Painter [66] and PromptDiffusion [69], both designed to handle multiple tasks using a unified framework, as state-of-the-art methods. We also report the results of other competing methods, which are specially trained on single tasks and do not use a general framework, for reference purposes. Due to limited computing resources, we cannot jointly train ICT on data from all tasks to achieve convergence in an affordable time. Therefore, we mainly report the results of single-task models (ICT-st) that are separately trained for each task, and single-category models (ICT-sc) that are jointly trained on data from multiple tasks of the same category. Nevertheless, we train a multi-category model (ICT-mc) on data from three tasks belonging to distinct categories to demonstrate our ICT’s validity in tackling various tasks using a single set of model parameters.
Implementation Details. We utilize the same training settings of SD to optimize ICT. We accumulate gradients every 16 batches with a batch size of 64. The learning rate is fixed to $6.4\times 10^{-5}$. All training images are resized to $256\times 256$ and we train ICT on 4 RTX 3090 GPUs.
4.2 Results and Analyses

### v2

@@ Abstract
We propose In-Context Translation (ICT), a general learning framework to unify visual recognition (e.g., semantic segmentation), low-level image processing (e.g., denoising), and conditional image generation (e.g., edge-to-image synthesis). Thanks to unification, ICT significantly reduces the inherent inductive bias that comes with designing models for specific tasks, and it maximizes mutual enhancement across similar tasks. However, the unification across a large number of tasks is non-trivial due to various data formats and training pipelines. To this end, ICT introduces two designs. Firstly, it standardizes input-output data of different tasks into RGB image pairs, e.g., semantic segmentation data pairs an RGB image with its segmentation mask in the same RGB format. This turns different tasks into a general translation task between two RGB images. Secondly, it standardizes the training of different tasks into a general in-context learning, where “in-context” means the input comprises an example input-output pair of the target task and a query image. The learning objective is to generate the “missing” data paired with the query. The implicit translation process is thus between the query and the generated image. In experiments, ICT unifies ten vision tasks and showcases impressive performance on their respective benchmarks. Notably, ICT performs well across three major categories of computer vision tasks, while its two competitors (Painter and PromptDiffusion) are only effective in at most two of these task categories. In addition, compared to its competitors, ICT trained on only 4 RTX 3090 GPUs is shown to be more efficient and less costly in training.
Index Terms:
In-Context learning, image translation, unified models.
I Introduction
@@
In the computer vision (CV) community, a unified framework for different tasks is also a long-standing aspiration. The unification is appealing because it side-steps task-specific designs, minimizes the inductive bias inherent in crafting models for different tasks, and maximizes the potential of mutual enhancement among related tasks. However, the progress of such unification in CV lags much behind NLP, due to two main Challenges. C1: NLP has the language token as an intuitive and unified data format for all tasks. However, CV encompasses highly heterogeneous data formats, e.g., images, semantic segmentation masks, keypoints, depth maps, etc. Such variation impedes the unification of expert models of different CV tasks. C2: Most NLP tasks can be easily converted into a sequence completion task [4] as language itself is sequential. In CV, the popular solution for image completion is Masked Autoencoders (MAE) [12]. The objective of MAE is to acquire an image representation by reconstructing the original image from a version where a random subset of image patches is masked out. This representation can be tailored to various specific discrimination tasks through fine-tuning. However, it is not intuitive how it can be applied to unify other categories of vision tasks, such as conditional image generation.
@@
Thanks to ICT’s unifying designs, in practice, it can be deployed to train various models according to the degree of unification. 1) When trained on joint data of different categories of tasks (e.g., semantic segmentation, image denoising, and depth-to-image synthesis), it yields a compact model that can generate outputs for every trained task. 2) When trained on joint data of a specific category (e.g., mask-to-image and depth-to-image data for the category of conditional image generation), it derives a single-category model. 3) When trained on data of an individual task, it learns a dedicated single-task model. This is a special case where the model itself is not unified. We highlight that the above three models are trained in different regimes and datasets, yielding distinct model parameters, but the training framework (including input-output format, backbone, training method, etc.) always remains the same, underscoring the “unifying” essence.
@@
Our experiments showcase extensive results of ICT in three types consisting of ten tasks in total. Our key observation is that ICT exhibits promising performance on all categories of tasks, while its competitors Painter [13] and PromptDiffusion [14] collapse on at least one of them (Fig. 1). Our contribution is thus three-fold:
  • A general learning framework that can cope with visual recognition, low-level image processing, and conditional image generation tasks.
  • A new in-context translation method that can be built on Stable Diffusion (SD), adapting SD’s pre-trained knowledge for downstream tasks in a unified fashion.
  • Extensive experiments on ten vision tasks across three categories and for three types of model training, by which we hope to spur more interesting research on how to induce a profound understanding of vision tasks through a unified learning framework.
II Related Works
II-A Unified Vision Models
Encouraged by the success of language generalist models [4, 5, 18], training a generalist model to unify different computer vision tasks has attracted significant interest in recent years. We identify three key components of designing a unified vision model: 1) data format, 2) backbone, and 3) training pipeline. Regarding these, we discuss related works and their primary differences from ours.
Some attempts [19, 20, 21, 22, 23, 24, 25] map the input image into discrete representations and implement sequence-to-sequence learning in the discrete space to handle different tasks. Two representative works, Unified IO [23] and Unified-IO 2 [24], homogenize various vision data modalities into a sequence of discrete vocabulary tokens and utilizes VQGAN [26] to support dense prediction tasks. This unification of data format via the discretization process leads to lossy data compression, making the solution suboptimal. In contrast, our unified RGB image format better suits vision tasks.
@@ II-A Unified Vision Models
Some recent works [30, 31, 32, 33] exploit the LLM-centric solution: they align different modalities into a unified space defined by a pre-trained LLM (e.g., Vicuna [34]) and invoke expert models to execute different tasks (e.g., Stable Diffusion [15] for text-to-image generation and SEEM [35] for image segmentation). However, these vision LLMs still rely on various task-specific models, which is highly inefficient when the number of tasks increases. In addition, sophisticated instruction-based data curation is needed for training.
Similar to our ICT, a bunch of works [36, 13, 14, 37, 38, 39] utilize the RGB image as a general data format to unify vision tasks. MAE-VQGAN [36] and Painter [13] propose a masked image generation framework, where the input image and an example pair are stitched together and the model needs to predict the masked region (similar to MAE [12]). They adopt a pre-trained ViT-based MAE as the backbone, limiting their performance in image generation. ICT does not suffer from this issue due to the generative backbone we use. In addition, they apply random masking of image patches during training while ICT is trained to predict the whole desired output image that is more in line with the inference scenario. ICT supports the usage of textual prompt to enable more controlled image generation, which is beyond the reach of MAE-VQGAN and Painter.
@@
InstructCV [37] and InstructDiffusion [38] learn a unified model based on a generative backbone and cast different vision tasks as text-guided image editing. Nevertheless, they heavily rely on delicate prompt design, limiting their applications to vision tasks that are difficult to instruct by using human language (e.g., InstructDiffusion cannot handle the task of depth estimation). ICT refrains from this issue by applying a simple image-based context to inform the desired task.
II-B Diffusion Models
Diffusion models [41, 42, 43, 44, 45] have recently become the primary choices for generative modeling of data. Following the definition in Denoising Diffusion Probabilistic Models (DDPM) [42], a diffusion model consists of a forward diffusion process that gradually adds noise to data and a reverse denoising process that reconstructs desired data from the noise. Recent methods based on diffusion models achieve state-of-the-art results in many vision tasks [46], including image/video generation [47, 48], image editing [49, 50], image-to-image translation [51], low-level image processing [52, 53], etc. Notably, large-scale text-to-image diffusion models show compelling ability [54, 55, 15, 56] to generate high-fidelity visual content. These pre-trained models are broadly utilized in applications such as concept customization [57, 58, 59] and image composition [60]. Some recent works [61, 62, 63, 64, 65] further reveal the potential of applying pre-trained diffusion models to discriminative tasks, which encourages us to build a general framework that unifies generative and discriminative model training.
II-C In-Context Learning
GPT-3 [4] has demonstrated the ability of pre-trained LLMs to adapt to various downstream NLP tasks given example pairs. After that, there have been efforts in exploring the ability of in-context learning [7, 8, 9, 66, 10, 67]. A prominent track of works [68, 69, 70, 71] fine-tunes language models on a collection of datasets described via instructions (e.g., task prompt, demonstration examples, and constraints), and attains significant improvement of the model’s generalization on unseen tasks. This in-context learning pipeline has recently been introduced to vision-language tasks as well [11, 72, 73, 74]. A representative work, Flamingo [11], bridges pre-trained vision and language models by fine-tuning them on text-image instruction-following data and showcases impressive few-shot results in a variety of tasks such as image captioning and visual question-answering. By comparison, ICT exploits a new approach of in-context translation, which is based on the use of an image-label pair as well as an optional prompt for both image- and text-level context.
III In-Context Translation (ICT)
@@
To learn function $f$, we first construct context-based training data by converting the input-output data into RGB images consistently across different vision tasks (Section III-A). Then, we build the ICT framework on top of SD [15] and offer different training strategies according to the unification degrees (Section III-B). For example, for the most unification, each model training iteration uses a batch of context-based data sampled from ten vision tasks (across three types).
III-A Context-based Data Construction
@@
Visual Recognition. We consider three representative recognition tasks, including semantic segmentation, depth estimation, and keypoint detection. 1) Semantic segmentation is a dense classification task wherein the input is an image and the output is per-pixel semantic labels. We transfer the output to an RGB image by assigning a color to each pixel according to a predefined semantic-color codebook, ensuring that each semantic class has its unique RGB value [13]. During inference, we obtain the predicted semantic class of each pixel by finding its “nearest” color in the codebook. 2) Depth estimation is a dense regression task that needs to predict the depth value of each pixel of the input image. To convert the output, i.e., one-channel depth map, into an RGB image, we linearly scale the range of pixel values to $\left[0,255\right]$ and replicate it three times to yield a three-channel “image”. 3) Keypoint detection aims to locate key object components on the input image. We formulate the output as a RGB image by first drawing squares (each square occupies 9$\times$9 pixels) whose center is the location of keypoints, and then rendering each square with a unique color associated with the key component.
@@
III-B Training Framework
We implement ICT to run an in-context image translation process on top of SD. SD is a text-to-image model that incorporates a diffusion process in the latent space of a pre-trained autoencoder. Specifically, a denoising U-Net is trained to fit the distribution of latent codes, which models the reverse diffusion process. Taking the noisy latent and the time step as input, this U-Net is further conditioned on the textual embeddings extracted by the text encoder of CLIP [78]. Visual and textual embeddings interact by cross-attention inside the U-Net to produce the output of the current time step. During inference, SD performs iterative reverse diffusion on a randomly sampled noise to generate an image that faithfully adheres to the input text. In our work, to fulfill (1) for various visual tasks, we fine-tune the U-Net on our context-based datasets.
TABLE I: Task and Dataset Configuration
Task Dataset Training images Testing images
Semantic segmentation ADE20K [79] 20,210 2,000
Depth estimation NYUv2 [80] 24,231 654
Keypoint detection COCO [81] 149,781 6,352
Image deraining Merged deraining datasets [82, 83, 84, 85, 86] 13,712 4,300
Image denoising SIDD [87] 96,000 1,280
Low-light image enhancement LoL [88] 485 15
Mask-to-image synthesis ADE20K [79] 20,210 2,000
Depth-to-image synthesis NYUv2 [80] 24,231 654
Pose-to-image synthesis COCO [81] 149,781 6,352
Edge-to-image synthesis COCO [81] 118,287 5,000
@@ III-B Training Framework
$\mathcal{L}(\theta)=\mathbb{E}_{z,m,b,y,\epsilon,t}\Bigl{[}\lVert\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}(m),b,c_{\phi}(y))\rVert_{2}^{2}\Bigr{]},$ (2)
where $z$ is the latent code extracted from the ground truth, $y$ is the input text, $\epsilon\sim\mathcal{N}(0,1)$ is a noise term, $t$ is the time step, $\epsilon_{\theta}$ is the denoising U-Net, $z_{t}$ is the noisy version of $z$ at time $t$, $\mathcal{E}$ is the VAE encoder, and $c_{\phi}$ is the text encoder. We fine-tune the U-Net and keep the text encoder and the autoencoder frozen.
@@
IV Experiments
IV-A Experimental Settings
TABLE II: Comparison Results on Visual Recognition. : Specialized Methods for Each Task. : Officially Trained Painter Model Using 32$\times$ the Computing Power of ICT. : Retrained Painter Using Its Official Code under the Same Computing Resources as ICT. §: A Concurrent Work (Results Are Obtained Using the Official Weights). Bold: Best. Underline: Second Best. We Ignore Specialized Models When Ranking Best and Second Best and This Applies to All Tables. The Results of ICT Are Reported as the Average Scores and Standard Deviations across Three Trials
@@ IV-A Experimental Settings
TABLE III: Comparison Results on Low-Level Image Processing. : Specialized Methods for Each Task. : Officially Trained Painter Model Using 32$\times$ the Computing Power of ICT. : Retrained Painter Using Its Official Code under the Same Computing Resources as ICT. : Following InstructDiffusion [38], It Directly Reconstructs the Ground Truth via the Autoencoder of Pre-trained SD, and the Corresponding Results Indicate the Upper Bound of ICT. Bold: Best. Underline: Second Best
@@
TABLE IV: Comparison Results on Conditional Image Generation. : Specialized Methods for Each Task. : Trained Using Official Code under the Same Computing Resources as ICT. Note That There Is No Officially Trained Painter Model for Conditional Image Generation. Bold: Best. Underline: Second Best
@@
ControlNet [76] 35.4 36.5 43.9 0.327 43.0 0.479 12.9 0.685
Painter [13] 75.7 23.2 89.3 0.417 200.1 0.709 233.1 0.271
PromptDiffusion [14] 31.0 30.5 52.5 0.528 40.6 0.457 13.8 0.690
ICT-st 29.9 $\pm$ 0.3 28.7 $\pm$ 0.2 44.0 $\pm$ 0.7 0.374 $\pm$ 0.005 34.7 $\pm$ 0.3 0.359 $\pm$ 0.009 13.6 $\pm$ 0.2 0.671 $\pm$ 0.005
ICT-sc 27.8 $\pm$ 0.6 28.5 $\pm$ 0.4 44.2 $\pm$ 0.8 0.372 $\pm$ 0.004 34.3 $\pm$ 0.5 0.430 $\pm$ 0.007 13.5 $\pm$ 0.4 0.673 $\pm$ 0.007
@@
Datasets. We conduct experiments involving ten vision tasks on six datasets, including ADE20K [79], NYUv2 [80], COCO [81], merged deraining datasets [82, 83, 84, 85, 86], SIDD [87], and LoL [88]. We adopt the same training/testing split as Painter [13]. Please refer to Table I for a detailed task and dataset configuration.
@@
Methods. We evaluate ICT with its two direct competitors, Painter [13] and PromptDiffusion [14], both designed to handle multiple tasks using a unified framework, as state-of-the-art methods. Some small-scale comparisons with related works including InstructCV [37], Unified-IO 2 [24], and InstructDiffusion [38] are conducted as well. We also report the results of other competing methods, which are specially trained on single tasks and do not use a general framework, for reference purposes. Due to limited computing resources, we cannot jointly train ICT on data from all tasks to achieve convergence in an affordable time. Therefore, we mainly report the results of single-task models (ICT-st) that are separately trained for each task, and single-category models (ICT-sc) that are jointly trained on data from multiple tasks of the same category. Nevertheless, we train a multi-category model (ICT-mc) on data from three tasks belonging to distinct categories to demonstrate our ICT’s validity in tackling various tasks using a single set of model parameters.
Implementation Details. We utilize the same training settings of SD to optimize ICT. We adopt a smooth L1 version of (2) to train our model. We accumulate gradients every 16 batches with a batch size of 64. The learning rate is fixed to $6.4\times 10^{-5}$. All training images are resized to $256\times 256$ and we train ICT on 4 RTX 3090 GPUs. During training, we randomly drop 10% text-conditioning to improve classifier-free guidance sampling [95]. We load pre-trained weights from Stable Diffusion-v1.5-inpainting for fine-tuning. We randomly sample an input-output pair as the image instruction during training and adopt a fixed pair from the training dataset (same as Painter [13]) as the image context at the inference time. By setting the random noise in the reverse diffusion process to 0 (i.e., a deterministic sampling), DDIM [43] manages to generate an image with fewer sampling steps compared to DDPM. We adopt DDIM sampling with 50 steps for inference.
IV-B Results and Analyses
Figure 3: Visual comparison results on visual recognition. “GT” represents ground truth. (a) PromptDiffusion [14] tends to mispredict pixels as “black”. Red arrow indicates the mispredictions produced by Painter [13]. (b) The competing methods, especially PromptDiffusion [14], fail to consistently produce accurate depth predictions (red arrow). (c) PromptDiffusion [14] tends to overpredict keypoints (red arrow). In contrast, the proposed ICT consistently produces accurate predictions.
Figure 4: Visual comparison results on low-level image processing. “GT” represents ground truth. PromptDiffusion [14] tends to generate distorted results or images with color shifts that do not adhere to the query image.
Figure 5: Visual comparison results on conditional image generation. “GT” represents ground truth. Painter [13] failed in generating realistic images from conditions with sparse semantics.
Visual Recognition Results. We assess the proposed ICT on three visual recognition tasks described in Section III-A. Standard metrics are adopted for evaluation: 1) mean Intersection-over-Union (mIoU) for semantic segmentation; 2) root mean squared error (RMSE), absolute relative error (REL), and the accuracy under the threshold ($\delta_{1}<1.25$) for depth estimation.